OpenStudy (anonymous):

FAN AND MEDAL! Suppose 10 people share 15 cinnamon rolls. Each person gets same amount. What fraction of a roll does each recieve in mixed number and fraction form?

OpenStudy (anonymous):

2/3 or .66666...

OpenStudy (anonymous):

... wut bout m.n?

OpenStudy (anonymous):

how am i supposed to answer?

OpenStudy (anonymous):

I mean 3/2 or 1 1/2

OpenStudy (anonymous):

\[\frac{ 15 }{ 10 }=1\frac{ 1 }{ 2 }=\frac{ 3 }{ 2}\]